# Randomized block-Krylov subspace methods for low-rank approximation of matrix functions

### v1

@@ 2.3 Krylov-aware low-rank approximation
A key observation in [6] was that $\operatorname{range}(\bm{W})\subseteq\operatorname{range}(\bm{Q}_{s})$, where $\bm{W}$ and $\bm{Q}_{s}$ are as in Algorithm 3. Therefore, by [33, Lemma 3.3] one has
@@
Since $\bm{Q}_{s}$ could have as many as $s\ell$ columns, an apparent downside to this approach is that approximating $f(\bm{A})\bm{Q}_{s}$ might require $rs\ell$ matvecs with $\bm{A}$ if we use $r$ iterations of the block Lanczos method (Algorithm 1) to approximately compute each matvec with $f(\bm{A}$. The key observation which allows Krylov-aware algorithms to be implemented efficiently is the following:
Lemma 2.2 ([6, Section 3]).
@@
In addition to the work of Chen and Hallman, this observation was recently used to analyze “small-block” randomized Krylov methods for low-rank approximation [27] and a randomized variant of the block conjugate gradient algorithm [7].
@@
where the last inequality is due to the fact that the spectrum of $\bm{T}_{q}$ is contained in $[\lambda_{\min},\lambda_{\max}]$. Optimizing over $p\in\mathbb{P}_{2r+1}$ gives the result. ∎
We can now present the Krylov-aware low-rank approximation algorithm; see Algorithm 4. The total number of matvecs with $\bm{A}$ is $(s+r)\ell$, the same as Algorithm 3. However, as noted above, Algorithm 4 (approximately) projects $f(\bm{A})$ onto a higher dimensional subbspace, ideally obtaining a better approximation.
@@
input: Symmetric $\bm{A}\in\mathbb{R}^{n\times n}$. Rank $k$. Oversampling parameter $\ell-k$. Matrix function $f:\mathbb{R}\to\mathbb{R}$. Number of iterations $q=s+r$.
@@
3:Compute $\bm{X}=f(\bm{T}_{q})_{1:d_{s,\ell},1:d_{s,\ell}}$. $\triangleright$ $\approx\bm{Q}_{s}^{\textsf{{T}}}f(\bm{A})\bm{Q}_{s}$
@@ 3.1 Error bounds for inexact projections
In this section we will derive error bounds for $\|f(\bm{A})-\bm{Q}\bm{X}\bm{Q}^{\textsf{{T}}}\|_{\textsf{{F}}}$ and $\|f(\bm{A})-\bm{Q}\llbracket\bm{X}\rrbracket_{k}\bm{Q}^{\textsf{{T}}}\|_{\textsf{{F}}}$ where $\bm{Q}$ is any orthonormal basis and $\bm{X}$ is any matrix. By [33, Lemma 3.3] we know that the optimal choice of $\bm{X}$ is $\bm{X}=\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}$. However, since $\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}$ can only be computed approximately, we need to show that the errors $\|f(\bm{A})-\bm{Q}\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}\bm{Q}\|_{\textsf{{F}}}$ and $\|f(\bm{A})-\bm{Q}\llbracket\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}\rrbracket_{k}\bm{Q}\|_{\textsf{{F}}}$ are robust against perturbations in $\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}$. Theorem 3.1 provides such a result.
@@
Given an orthonormal basis $\bm{Q}$ and an approximation $\bm{X}$ to $\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}$. Then,
$\|f(\bm{A})-\bm{Q}\bm{X}\bm{Q}^{\textsf{{T}}}\|_{\textsf{{F}}}^{2}=\|f(\bm{A})-\bm{Q}\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}\bm{Q}^{\textsf{{T}}}\|_{\textsf{{F}}}^{2}+\|\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}-\bm{X}\|_{\textsf{{F}}}^{2},$ (6)
and
$\|f(\bm{A})-\bm{Q}\llbracket\bm{X}\rrbracket_{k}\bm{Q}^{\textsf{{T}}}\|_{\textsf{{F}}}\leq\|f(\bm{A})-\bm{Q}\llbracket\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}\rrbracket_{k}\bm{Q}^{\textsf{{T}}}\|_{\textsf{{F}}}+2\|\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}-\bm{X}\|_{\textsf{{F}}}.$ (7)
@@
(6) is immediate due to the Pythagorean theorem.
We now proceed with proving (7) using a similar argument to [41, Proof of Theorem 5.1]. Define $\bm{C}=f(\bm{A})-\bm{Q}\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}\bm{Q}^{\textsf{{T}}}+\bm{Q}\bm{X}\bm{Q}^{\textsf{{T}}}$. Note that $\|\bm{C}-f(\bm{A})\|_{\textsf{{F}}}=\|\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}-\bm{X}\|_{\textsf{{F}}}$ and $\bm{Q}^{\textsf{{T}}}\bm{C}\bm{Q}=\bm{X}$. Hence,
@@
Since $\bm{Q}\llbracket\bm{Q}^{\textsf{{T}}}\bm{C}\bm{Q}\rrbracket_{k}\bm{Q}^{\textsf{{T}}}$ is the best rank $k$ approximation whose range is contained in $\operatorname{range}(\bm{Q})$ [33, Lemma 3.3] (a similar result is given in [19, Theorem 3.5]), we have
@@
where we used that $\bm{W}\llbracket\bm{W}^{\textsf{{T}}}f(\bm{A})\bm{W}\rrbracket_{k}\bm{W}^{\textsf{{T}}}$ is the best rank $k$ approximation to $f(\bm{A})$ whose range and co-range is contained in $\operatorname{range}(\bm{W})$ [33, Lemma 3.3]. Hence, by Theorem 3.1 we have
@@ 3.2 Structural bounds
which quantifies the extent to which an $(s-1)$ degree polynomial can be large (relative to $f$) on the eigenvalues $\lambda_{1},\ldots,\lambda_{k}$ and small on the remaining eigenvalues. Recall that we order $\bm{A}$’s eigenvalues with respect to $f$, so $|f(\lambda_{1})|\geq|f(\lambda_{1})|\geq\ldots\geq|f(\lambda_{n})|$.
@@
Consider $\bm{A}\in\mathbb{R}^{n\times n}$ as defined in (1). Assuming $\bm{\Omega}_{k}$ in (2) has rank $k$, for all functions $f:\mathbb{R}\to\mathbb{R}$, we have
@@
From this bound, we can further see why Algorithm 4 should be preferred over directly applying the randomized SVD to $f(\bm{A})$. As an extreme case, consider when $f$ is a polynomial of degree at most $s-1$. Then the standard randomized SVD bound [21, Theorem 9.1] essentially replaces $p$ with $f$ in $\mathcal{E}_{\bm{\Omega}}(s;f)$ in (11):
$\|f(\bm{A})-\bm{W}\llbracket\bm{W}^{T}f(\bm{A})\bm{W}\rrbracket_{k}\bm{W}\|_{\textsf{{F}}}^{2}\leq\|f(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}^{2}+5\|f(\bm{\Lambda}_{n\setminus k})\bm{\Omega}_{n\setminus k}\bm{\Omega}_{k}^{\dagger}\|_{\textsf{{F}}}^{2},$
where $\bm{W}$ is an orthonormal basis for $\operatorname{range}(f(\bm{A})\bm{\Omega})$. Since $\mathcal{E}_{\bm{\Omega}}(s;f)$ minimizes over all degree $s-1$ polynomials, it is always smaller than $\|f(\bm{\Lambda}_{n\setminus k})\bm{\Omega}_{n\setminus k}\bm{\Omega}_{k}^{\dagger}\|_{\textsf{{F}}}^{2}$. So, the approximation returned by Algorithm 4 is expected to be more accurate compared to the randomized SVD. When $f$ is not a polynomial, the error of the randomized SVD roughly corresponds to plugging a good polynomial approximation for $f$ into Eq. 10, but again there might be a better choice to minimize $\mathcal{E}_{\bm{\Omega}}(s;f)$.
Indeed, in many cases, we find that the improvement obtained from effectively optimizing over all possible degree $s-1$ polynomials is significant – there is often a much better choice of polynomial than a direct approximation to $f$ for Eq. 10. This is reflected in our experiments (Section 4) and we also provide a concrete analysis involving the matrix exponential in Section 3.5.
@@
The first inequality is due to the fact that $\bm{Q}_{s}\bm{Q}_{s}^{\textsf{{T}}}f(\bm{A})\bm{Q}_{s}\bm{Q}_{s}^{\textsf{{T}}}$ is the nearest matrix to $f(\bm{A})$ in the Frobenius norm whose range and co-range is contained in $\operatorname{range}(\bm{Q}_{s})$ [33, Lemma 3.3]. We proceed with proving the second inequality.
To do so, it of course suffices to prove the inequality where $\mathcal{E}_{\bm{\Omega}}(s;f)$ is replaced with $\|p(\bm{\Lambda}_{n\setminus k})\bm{\Omega}_{n\setminus k}\bm{\Omega}_{k}^{\dagger}\|_{\textsf{{F}}}^{2}\cdot\max_{i=1,\ldots,k}\left|{f(\lambda_{i})}/{p(\lambda_{i})}\right|^{2}$ for any choice of $p\in\mathbb{P}_{s-1}$. Note that if we choose $p$ so that $p(\lambda_{i})=0$ for some $i=1,\ldots,k$ then the right hand side of (11) is infinite and the bound trivially holds. Hence, we may assume that $p(\lambda_{i})\neq 0$ for $i=1,\ldots,k$. Consequently, $p(\bm{\Lambda}_{k})$ is non-singular. Define $\bm{Z}=p(\bm{A})\bm{\Omega}\bm{\Omega}_{k}^{\dagger}p(\bm{\Lambda}_{k})^{-1}$ and let $\widetilde{\bm{P}}$ be the orthogonal projector onto $\operatorname{range}(\bm{Z})\subseteq\operatorname{range}(\bm{Q}_{s})$. Note that $\operatorname{rank}(\bm{Z})\leq k$ and $\bm{Q}_{s}\llbracket\bm{Q}_{s}^{\textsf{{T}}}f(\bm{A})\bm{Q}_{s}\rrbracket_{k}\bm{Q}_{s}^{\textsf{{T}}}$ is the best rank $k$ approximation to $f(\bm{A})$ whose range and co-range is contained in $\operatorname{range}(\bm{Q}_{s})$ [33, Lemma 3.3]. Hence,
@@
We are going to bound the two terms on the right hand side of (12) separately. Our analysis for the first is similar to the proof of [21, Theorem 9.1].
@@
where the inequality is due to [21, Proposition 8.2]. Consequently,
@@
We proceed with bounding the second term in (12). Our analysis is similar to the proof of [32, Lemma 3.7]. By the triangle inequality we have
@@
$\left\|\begin{bmatrix}\bm{0}&\\ &f(\bm{\Lambda}_{n\setminus k})\end{bmatrix}\widehat{\bm{P}}\right\|_{\textsf{{F}}}^{2}$ $=\operatorname{tr}\left(\begin{bmatrix}\bm{0}&\\ &f(\bm{\Lambda}_{n\setminus k})\end{bmatrix}\widehat{\bm{P}}\begin{bmatrix}\bm{0}&\\ &f(\bm{\Lambda}_{n\setminus k})\end{bmatrix}\right)$

### v2

@@ 2.3 Krylov-aware low-rank approximation
A key observation in [chen_hallman_23] was that $\operatorname{range}(\bm{W})\subseteq\operatorname{range}(\bm{Q}_{s})$, where $\bm{W}$ and $\bm{Q}_{s}$ are as in Algorithm 3. Therefore, by [funnystrom2, Lemma 3.3] one has
@@
Since $\bm{Q}_{s}$ could have as many as $s\ell$ columns, an apparent downside to this approach is that approximating $f(\bm{A})\bm{Q}_{s}$ might require $rs\ell$ matvecs with $\bm{A}$ if we use $r$ iterations of the block-Lanczos method (Algorithm 1) to approximately compute each matvec with $f(\bm{A})$. The key observation which allows Krylov-aware algorithms to be implemented efficiently is the following:
Lemma 2.2 (see e.g. [chen_hallman_23, Section 3]).
@@
In addition to the work of Chen and Hallman, this observation was recently used to analyze randomized Krylov methods for low-rank approximation [MeyerMuscoMusco:2024, moderateblocksize] and as an algorithmic tool in a randomized variant of the block conjugate gradient algorithm [BCGPreconditioning] and implementations of low-rank approximation algorithms [tropp2023randomized].
@@
where the last inequality is due to the fact that the spectrum of $\bm{T}_{q}$ is contained in $\lambda_{\min},\lambda_{\max}]$. Optimizing over $p\in\mathbb{P}_{2r+1}$ gives the result. ∎
We can now present the Krylov-aware low-rank approximation algorithm, which was presented in [chen_hallman_23, Section 3]; see Algorithm 4. The total number of matvecs with $\bm{A}$ is $(s+r)\ell$, the same as Algorithm 3.555Note that Algorithm 4 requires computing a matrix-function of size $(s+r)\ell\times(s+r)\ell$, which is more expensive than computing the two matrix functions of size $s\ell\times s\ell$ and $r\ell\times r\ell$ in Algorithm 3. However, if $s,r\ll n$, the cost of these matrix-function evaluations is not the dominant part of the computation. However, as noted above, Algorithm 4 (approximately) projects $f(\bm{A})$ onto a higher dimensional subspace, ideally obtaining a better approximation.
@@
input: Symmetric $\bm{A}\in\mathbb{R}^{n\times n}$. Rank $k$. Block-size $\ell\geq k$. Matrix function $f:\mathbb{R}\to\mathbb{R}$. Number of iterations $q=s+r$.
@@
3:Compute $\bm{X}=f(\bm{T}_{q})_{1:d_{s,\ell},1:d_{s,\ell}}$ where $d_{s,\ell}=\dim(\mathcal{K}_{s}(\bm{A},\bm{\Omega}))$. $\triangleright$ $\approx\bm{Q}_{s}^{\textsf{{T}}}f(\bm{A})\bm{Q}_{s}$
@@
Algorithm 4 requires $\ell(s+r)$ matvecs with $\bm{A}$ and $O(n(s+r)^{2}\ell^{2})$ additional operations, assuming full reorthogonalization is performed in Algorithm 1. For dense matrices, the cost is dominated by matvecs with $\bm{A}$, resulting in an overall complexity of $O(n^{2}(s+r)\ell)$ operations. In contrast, when $\bm{A}$ is sparse, the reorthogonalization becomes the dominant computational cost and scales quadratically with the block-size $\ell$. While Algorithm 2 and Algorithm 3 require the block-size $\ell$ to be larger than $k$ to be able to return a rank-$k$ approximation, Algorithm 4 can return a rank-$k$ approximation even if $\ell<k$, provided $q\ell\geq k$. Recent work has shown that single-vector methods (i.e., setting $\ell=1$) can often produce a more efficient low-rank approximation at the same computational cost. Furthermore, Chen and Hallman [chen_hallman_23] recommended using a small block-size, say $\ell=1,2,$ or $4$. Later in this work we will consider a version of Algorithm 4 which uses a single-vector starting block, i.e. $\ell=1$. However, its analysis relies on the analysis of the block-version in Algorithm 4, which restricts $\ell\geq k$. We therefore defer the discussion on a single-vector version of Algorithm 4 to Section 4. Moreover, while single-vector methods often produce more accurate approximations at the same number of arithmetic operations, large-block methods come with other advantages, such as being more parallelizable.
@@
Remark.
Modern methods for eigenpair approximation often employ restarting techniques to reduce memory requirements. Furthermore, Chen and Hallman does discuss a restarting strategy [chen_hallman_23, Section 4.2] based on classical techniques [sorensen, baglama]. While restarting techniques are well studied in the context of eigenpair approximations, they are less understood in the context of low-rank approximation. In this work, the theoretical analysis will not focus on restarting techniques.
@@ 3.1 Error bounds for inexact projections
In this section we will derive error bounds for $\|f(\bm{A})-\bm{Q}\bm{X}\bm{Q}^{\textsf{{T}}}\|_{\textsf{{F}}}$ and $\|f(\bm{A})-\bm{Q}\llbracket\bm{X}\rrbracket_{k}\bm{Q}^{\textsf{{T}}}\|_{\textsf{{F}}}$ where $\bm{Q}$ is any orthonormal basis and $\bm{X}$ is any matrix. By [funnystrom2, Lemma 3.3] we know that the optimal choice of $\bm{X}$ is $\bm{X}=\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}$. However, since $\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}$ can only be computed approximately, we need to show that the errors $\|f(\bm{A})-\bm{Q}\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}\bm{Q}\|_{\textsf{{F}}}$ and $\|f(\bm{A})-\bm{Q}\llbracket\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}\rrbracket_{k}\bm{Q}\|_{\textsf{{F}}}$ are robust against perturbations in $\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}$. Robustness of the former follows directly from the Pythagorean theorem
$\|f(\bm{A})-\bm{Q}\bm{X}\bm{Q}^{\textsf{{T}}}\|_{\textsf{{F}}}^{2}=\|f(\bm{A})-\bm{Q}\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}\bm{Q}^{\textsf{{T}}}\|_{\textsf{{F}}}^{2}+\|\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}-\bm{X}\|_{\textsf{{F}}}^{2}.$ (6)
Establishing robustness of the latter requires more care, especially since rank-$k$ truncations are not Lipschitz continuous. However, we show in Theorem 3.1 that it is indeed robust.
@@
Given an orthonormal basis $\bm{Q}\in\mathbb{R}^{n\times d}$ and an approximation $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{X}\in\mathbb{R}^{d\times d}$} to $\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}$. Then,
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\|f(\bm{A})-\bm{Q}\llbracket\bm{X}\rrbracket_{k}\bm{Q}^{\textsf{{T}}}\|_{\textsf{{F}}}\leq\|f(\bm{A})-\bm{Q}\llbracket\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}\rrbracket_{k}\bm{Q}^{\textsf{{T}}}\|_{\textsf{{F}}}+2\|\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}-\bm{X}\|_{\textsf{{F}}}.$} (7)
@@
We will prove (7) using a similar argument to [tropp2016randomized, Proof of Theorem 5.1]. Define $\bm{C}=f(\bm{A})-\bm{Q}\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}\bm{Q}^{\textsf{{T}}}+\bm{Q}\bm{X}\bm{Q}^{\textsf{{T}}}$. Note that $\|\bm{C}-f(\bm{A})\|_{\textsf{{F}}}=\|\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}-\bm{X}\|_{\textsf{{F}}}$ and $\bm{Q}^{\textsf{{T}}}\bm{C}\bm{Q}=\bm{X}$. Hence,
@@
Since $\bm{Q}\llbracket\bm{Q}^{\textsf{{T}}}\bm{C}\bm{Q}\rrbracket_{k}\bm{Q}^{\textsf{{T}}}$ is the best rank $k$ approximation whose range is contained in $\operatorname{range}(\bm{Q})$ [funnystrom2, Lemma 3.3] (a similar result is given in [gu_subspace, Theorem 3.5]), we have
@@
where we used that $\bm{W}\llbracket\bm{W}^{\textsf{{T}}}f(\bm{A})\bm{W}\rrbracket_{k}\bm{W}^{\textsf{{T}}}$ is the best rank $k$ approximation to $f(\bm{A})$ whose range and co-range is contained in $\operatorname{range}(\bm{W})$ [funnystrom2, Lemma 3.3]. Hence, by Theorem 3.1 we have
@@ 3.2 Structural bounds
which quantifies the extent to which an (s1)(s-1) degree polynomial can be large (relative to $f$) on the eigenvalues $\lambda_{1},\ldots,\lambda_{k}$ and small on the remaining eigenvalues; similar quantities have appeared in other contexts [RUHE1984391, guttelphd]. Recall that we order $\bm{A}$’s eigenvalues with respect to $f$, so $|f(\lambda_{1})|\geq|f(\lambda_{1})|\geq\ldots\geq|f(\lambda_{n})|$.
@@
Consider $\bm{A}\in\mathbb{R}^{n\times n}$ as defined in (1) and let $\bm{Q}_{s}$ be an orthonormal basis for $\mathcal{K}_{s}(\bm{A},\bm{\Omega})$. Assuming $\bm{\Omega}_{k}$ in (2) has rank $k$, for all functions $f:\mathbb{R}\to\mathbb{R}$, we have
@@
From this bound, we can further see why Algorithm 4 should be preferred over directly applying the randomized SVD to $f(\bm{A})$. As an extreme case, consider when $f$ is a polynomial of degree at most s1s-1. Then the standard randomized SVD bound [rsvd, Theorem 9.1] essentially replaces $p$ with $f$ in $\mathcal{E}_{\bm{\Omega}}(s;f)$ in (11):
$\|f(\bm{A})-\bm{W}\llbracket\bm{W}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textsf{{T}}}}f(\bm{A})\bm{W}\rrbracket_{k}\bm{W}\|_{\textsf{{F}}}^{2}\leq\|f(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}^{2}+5\|f(\bm{\Lambda}_{n\setminus k})\bm{\Omega}_{n\setminus k}\bm{\Omega}_{k}^{\dagger}\|_{\textsf{{F}}}^{2},$
where $\bm{W}$ is an orthonormal basis for $\operatorname{range}(f(\bm{A})\bm{\Omega})$. Since $\mathcal{E}_{\bm{\Omega}}(s;f)$ minimizes over all degree s1s-1 polynomials, it is always smaller than $\|f(\bm{\Lambda}_{n\setminus k})\bm{\Omega}_{n\setminus k}\bm{\Omega}_{k}^{\dagger}\|_{\textsf{{F}}}^{2}$. So, the approximation returned by Algorithm 4 is expected to be more accurate compared to the randomized SVD. When $f$ is not a polynomial, the error of the randomized SVD roughly corresponds to plugging a good polynomial approximation for $f$ into Eq. 10, but again there might be a better choice to minimize $\mathcal{E}_{\bm{\Omega}}(s;f)$.
Indeed, in many cases, we find that the improvement obtained from effectively optimizing over all possible degree s1s-1 polynomials is significant – there is often a much better choice of polynomial than a direct approximation to $f$ for Eq. 10. This is reflected in our experiments (Section 5) and we also provide a concrete analysis involving the matrix exponential in Section 3.5.
We also highlight that this bound shows that if there is a polynomial $p\in\mathbb{P}_{s-1}$ that is large on $\lambda_{1},\ldots,\lambda_{k}$ and small on $\lambda_{k+1},\ldots,\lambda_{n}$, then one can construct a near-optimal rank $k$ approximation whose range is contained in $\operatorname{range}(\bm{Q}_{s})$; similar ideas have been used in the context of low-rank approximations of $\bm{A}$ [MM15, tropp2023randomized] and eigenvalue approximation [paige1971computation, saad1980rates, underwood1975iterative]. In particular, we expect that $\mathcal{E}(s;f)$ will decay rapidly as $s$ increases when $f$ is monotonically increasing or decreasing, since in these cases one can find explicit polynomials to upper bound $\mathcal{E}_{\bm{\Omega}}(s;f)$. For example, when $f$ is monotonic one can choose the polynomial to be a scaled and shifted Chebyshev polynomial.
@@
The first inequality is due to the fact that $\bm{Q}_{s}\bm{Q}_{s}^{\textsf{{T}}}f(\bm{A})\bm{Q}_{s}\bm{Q}_{s}^{\textsf{{T}}}$ is the nearest matrix to $f(\bm{A})$ in the Frobenius norm whose range and co-range is contained in $\operatorname{range}(\bm{Q}_{s})$ [funnystrom2, Lemma 3.3]. We proceed with proving the second inequality.
To do so, it of course suffices to prove the inequality where $\mathcal{E}_{\bm{\Omega}}(s;f)$ is replaced with $\|p(\bm{\Lambda}_{n\setminus k})\bm{\Omega}_{n\setminus k}\bm{\Omega}_{k}^{\dagger}\|_{\textsf{{F}}}^{2}\cdot\max_{i=1,\ldots,k}\left|{f(\lambda_{i})}/{p(\lambda_{i})}\right|^{2}$ for any choice of $p\in\mathbb{P}_{s-1}$. Note that if we choose $p$ so that $p(\lambda_{i})=0$ for some $i=1,\ldots,k$ then the right hand side of (11) is infinite and the bound trivially holds. Hence, we may assume that $p(\lambda_{i})\neq 0$ for $i=1,\ldots,k$. Consequently, $p(\bm{\Lambda}_{k})$ is non-singular. Define $\bm{Z}=p(\bm{A})\bm{\Omega}\bm{\Omega}_{k}^{\dagger}p(\bm{\Lambda}_{k})^{-1}$ and let $\widetilde{\bm{P}}$ be the orthogonal projector onto $\operatorname{range}(\bm{Z})\subseteq\operatorname{range}(\bm{Q}_{s})$. Note that $\operatorname{rank}(\bm{Z})\leq k$ and $\bm{Q}_{s}\llbracket\bm{Q}_{s}^{\textsf{{T}}}f(\bm{A})\bm{Q}_{s}\rrbracket_{k}\bm{Q}_{s}^{\textsf{{T}}}$ is the best rank $k$ approximation to $f(\bm{A})$ whose range and co-range is contained in $\operatorname{range}(\bm{Q}_{s})$ [funnystrom2, Lemma 3.3]. Hence,
@@
We are going to bound the two terms on the right hand side of (12) separately. Our analysis for the first is similar to the proof of [rsvd, Theorem 9.1].
@@
where $\preceq$ denotes the Loewner order [matrixanalysis, Definition 7.7.1] and the inequality is due to [rsvd, Proposition 8.2]. Hence, we have $f(\bm{\Lambda}_{1})\left(\bm{I}-(\bm{I}+\bm{F}^{\textsf{{T}}}\bm{F})^{-1}\right)f(\bm{\Lambda}_{1})\preceq f(\bm{\Lambda}_{1})\bm{F}^{\textsf{{T}}}\bm{F}f(\bm{\Lambda}_{1})$ and $f(\bm{\Lambda}_{2})\left(\bm{I}-\bm{F}(\bm{I}+\bm{F}^{\textsf{{T}}}\bm{F})^{-1}\bm{F}^{\textsf{{T}}}\right)f(\bm{\Lambda}_{2})\preceq f(\bm{\Lambda}_{2})^{2}$. Consequently, using [matrixanalysis, Corollary 7.7.4], the definition of $\bm{F}$, and strong submultiplicativity of the Frobenius norm we get
@@
We proceed with bounding the second term in (12). Our analysis is similar to the proof of [persson_kressner_23, Lemma 3.7]. By the triangle inequality we have
@@
[…]